### v1

# Zero-Shot Position Debiasing for Large Language Models

Abstract
[…]
1 Introduction
[…]
LLM have demonstrated remarkable unsupervised ability in various tasks (Kojima et al., 2022). Even so, fine-tuning still overtakes it under the task-specific setting (Ding et al., 2023). However, fine-tuned LLMs might rely on the dataset biases and artifacts as shortcuts for prediction, as the fine-tuning datasets are sometimes skewed due to budget constraints (Du et al., 2022). Consequently, these LLMs have poor generalization performance when applied to unseen test data and are vulnerable to various types of adversarial attacks (Meade et al., 2022).
Position bias has been shown in LLMs after fine-tuning (Liu et al., 2023). Specially, the well-known LLMs, e.g., GPT-3.5111https://openai.com/blog/gpt-3-5-turbo-fine-tuning-and-api-updates, longchat-13B [^2] https://lmsys.org/blog/2023-06-29-longchat, are skilled when the relevant information occurs at the beginning or end of the input context, while the performance significantly degrades when LLMs need to find relevant information in the middle of the context. Analysis of conversational question answering (CQA) on CANARD (Elgohary et al., 2019) dataset further confirms the existence of position bias. As shown in Fig. 1, after fine-tuning, T5-FT achieves about 0.8 in terms of ROUGE-L when relative position is 1, but less than 0.1 when relative position is 4. The striking contrast in performance attracts increasing attention from researchers on debiasing or mitigating position bias Meade et al. (2022).
Figure 1: Question answering performance of T5-large (T5) and fine-tuned T5-large (FT) over different relative positions in CANARD. Relative position means the distance between utterances containing the last turn answer span and the current turn answer span.
Early works mainly focus on mitigating position bias on extractive question answering (QA) before the emergence of LLMs (Ko et al., 2020; Karimi Mahabadi et al., 2020). A prominent debiasing method is Product-of-Expert (PoE), which discourages the extractive model from learning position bias picked up by the fixed biased model (Du et al., 2021; Shinoda et al., 2022). Recently, more and more works focus on generative language model debiasing for LLMs (Guo et al., 2022; Li et al., 2023). These works mostly focus on mitigating social bias (Kasneci et al., 2023), e.g., gender bias, and racial bias. They adopt either in-context learning (ICL) to guide the generation of LLM (Meade et al., 2023) or prompt tuning (PT) to fine-tune prompts for LLMs (Li et al., 2023). However, these works cannot be transferred to mitigate position bias, as they require manually annotated non-biased samples for ICL or external bias knowledge for PT, which are lacking for position bias.
In this paper, we propose to mitigate position bias for LLMs in zero-shot setting, i.e., without external bias knowledge or any non-biased samples. To achieve this, we propose to leverage the low position bias characteristics of pre-trained LLMs. Previous studies have shown that pre-trained LLMs are more robust to position bias (Utama et al., 2021). After being pre-trained on a large volume of data, pre-trained LLMs can leverage utterances from various positions in the document for generation, thus having low position bias. As shown in Fig. 1, pre-trained T5 achieves ROUGE-L score (Lin, 2004) ranging from 0.2 to 0.4 across almost all relative positions, demonstrating its robustness against position bias.
Motivated by this, we propose a ZOE framework to mitigate position bias for LLMs. First, we use a low-bias inference module to collect unsupervised responses with low position bias by applying diverse prompting strategies. Then, we propose a MSA module to prune the unsupervised responses, as low-quality responses will also undermine model performance on non-biased samples. Finally, we use a multi-objective optimization module to additionally optimize with unsupervised responses. The whole process does not require any external bias knowledge or non-biased samples, which is zero-shot and general enough.
To verify the effectiveness of ZOE, we have done experiments on five tasks with eight datasets. Experimental results show that the ZOE achieves superior performance in mitigating three types of position bias significantly, including lead bias, relative position bias and lexical bias. To summarize,
  • We propose to mitigate position bias in a zero-shot setting, i.e., without any external knowledge or datasets.
  • We propose a ZOE framework with a MSA module to prune low-quality unsupervised responses for fine-tuning.
  • Experiments show that ZOE mitigates various types of position biases by sacrificing only small performance on biased samples, demonstrating its effectiveness and generalization.
2 Preliminary
2.1 Task Definition
Given a biased dataset $D$ for the target task, the task aims to fine-tune the pre-trained model to improve the robustness against position bias, i.e., to achieve superior performance on non-biased samples by retaining the performance on biased samples. Here, the biased samples exhibit position bias consistent with that of the training samples and the non-biased samples lie outside the distribution of the training samples. As the positions of useful utterances vary in real-world scenarios, it is more applicable to achieve improved performance on non-biased samples by sacrificing partial performance on biased samples.
[…]
Figure 2: Overview of our proposed zero-shot position debiasing (ZOE) framework (taking CQG as the example). It consists of three components: low-bias inference to collect unsupervised responses, master-slave alignment module to align unsupervised responses, multi-objective optimization to fine-tune with unsupervised responses.
2.2 Large Language Model
[…]
LLM have attracted much attention and become state-of-the-art due to their remarkable ability of language generation. They formulate all natural language processing (NLP) tasks as language generation tasks with different task prompts:
$\begin{split}p(y)&=p(y|prompt,x)\\ &=\prod_{t}p(y_{t+1}|prompt,x,y_{1},y_{2},...,y_{t}),\end{split}$ (1)
where $x$, $y$ and $prompt$ are task input, output, and task prompt, respectively. And $y_{i}$ means the $i$-th token in $y$. Task prompt $prompt$ consists of the task instruction and demonstrations to tell the LLMs the definition of the task and how it works. The introduction of task prompt enables LLMs to utilize all available data for training and improve their generalization ability on unseen tasks.
3 Model
We propose a zero-shot position debiasing (ZOE) framework to mitigate position bias for generative LLMs. As shown in Fig. 2, ZOE consists of three parts: low-bias inference, master-slave alignment (MSA) and multi-objective optimization, where all modules do not require external bias knowledge or non-biased dataset. Low-bias inference generates unsupervised responses with lower position bias based on pre-trained LLMs (in §3.1). Then we prune these low-quality unsupervised responses based on the target responses by master-slave alignment (in §3.2). Finally, we fine-tune the LLMs by optimizing two objectives: target responses and aligned unsupervised responses (in §3.3).
3.1 Low-bias Inference
Given a biased training dataset $D=\{(x_{i},y_{i})\}_{i=1}^{N}$ for the target task, the low-bias inference module aims to generate unsupervised low-bias responses based on pre-trained LLMs, which are more robust and have low position bias Utama et al. (2021).
We generate unsupervised responses through three prompting strategies and apply different prompting strategies based on the task type.
  • Instruction-only prompting generate responses of target task by feeding the task input and task instruction to the pre-trained LLMs. Concretely, we assign the $prompt$ by task instruction in Eq. 1 for the generation. We adopt this prompting strategy for three fact generation tasks, CQA, KGC and summarization, as their target responses are always unique and unambiguous in semantics.
[…]
  • Diverse prompting obtains multiple responses of the target task by feeding various prompts to LLMs. It can be used in CQG, as various prompts can specify diverse aspects of the generated responses. For instance, we can specify a ‘What’ question for generation by feeding “Please generate a ‘What’ question based on the following dialogue context and passage”.
3.2 Master-slave Alignment
Given the generated unsupervised response $y^{\prime}_{i}$ for input $x_{i}$, and the target response $y_{i}$, the master-slave alignment (MSA) module prunes the unsupervised responses to align with the target response. The reason is that unsupervised responses are low-quality, as they are generated from pre-trained LLMs without training on the target task, which only serves as a slave objective for mitigating position bias. In contrast, the target responses are human-annotated with high quality, which is the primary or master objective for optimization.
Alignment for NLI. As the generated responses for NLI are deterministic, i.e., entailment, neutral and contradiction, we first estimate the probability distribution over all classes by prompting and then align it by masking. As maximizing the probability of the target class will strengthen the position bias of the training dataset, we align the probability distribution by masking the target class to mitigate position bias:
$y^{\prime}_{i}=s_{i}\cdot mask_{i},$ (2)
where $mask_{i}$ is a vector to mask the target class:
$mask_{i,j}=\begin{cases}0,\text{ if }class_{j}=y_{i}\\ 1,\text{otherwise}.\end{cases}$ (3)
Here, $class_{j}$ is the tokens for $j$-th class and $y_{i}$ is the tokens for the target class. $s_{i}$ is the probabilities distribution over all classes in NLI inference:
$\begin{split}s_{i}&=[s_{i,1},s_{i,2},...,s_{i,|class|}]\\ s_{i,j}&=\frac{p(class_{j}|x_{i})}{\sum_{j}^{|class|}p(class_{j}|x_{i})},\end{split}$ (4)
where $p(class_{j}|x_{i})$ is the generation probabilities of tokens for $j$-th class, which is calculated by Eq. 1.
Alignment for other tasks. We first identify low-quality unsupervised responses and then drop them, as modifying after identifying responses will lead to error accumulation. Four types of low-quality responses are identified in CQA, summarization, KGC and CQG: non-compliant responses, dull responses, incoherent responses and unreliable responses. We apply and combine four strategies to identify them.
  • Non-compliant identification identifies whether the unsupervised responses follow the task instruction by keyword matching. For example, it checks whether the generated question is a ‘What’ question when specified in the instruction.
  • Dull identification identifies whether the unsupervised responses are dull by keyword matching, e.g., “What is the title of the passage?”.
  • Incoherent identification identifies whether the unsupervised responses are incoherent based on perplexity. The response is dropped if the perplexity of any token in it falls below a pre-defined threshold.
  • Unreliable identification identifies whether the unsupervised responses are reliable by comparing them with the target response. The intuition is that the fact in the response may be wrong if its semantics deviate significantly from the fact in the reference. We identify and drop the response if the overlap score between it and the target response is less than a pre-defined threshold.
For CQA, summarization and KGC tasks, we align unreliable responses as the facts in answers, summaries and knowledge-enriched responses are always unique in semantics. For CQG task, we align non-compliant, dull or incoherent responses as the appropriate questions are diverse in semantics.
3.3 Multi-Objective Optimization
Given the target response $y_{i}$, aligned unsupervised response $y^{\prime}_{i}$ and input $x_{i}$, the multi-objective optimization module fine-tunes the model to generate high-quality but low-bias responses. It fine-tunes the model by optimizing two objectives: target responses as a master objective to improve the generation quality and aligned unsupervised responses as a slave objective to mitigate position bias:
$\begin{split}\mathcal{L}=&(1-\alpha)\cdot\mathcal{L}_{target}(x_{i},y_{i})+\\ &\alpha\cdot\mathcal{L}_{align}(x_{i},y^{\prime}_{i}),\end{split}$ (5)
where $\alpha$ is a hyper-parameter and $\mathcal{L}_{target}(x_{i},y_{i})$ is a negative log-likelihood (NLL) loss to maximize the generation probability of target response:
$\mathcal{L}_{target}(x_{i},y_{i})=-\sum_{j=1}^{|y_{i}|}\log p(y_{i,j}|x_{i},y_{i,<j}).$ (6)
$\mathcal{L}_{align}(x_{i},y^{\prime}_{i})$ is a task-specific objective to mitigate position bias.
For NLI task, we maximize the probability of the most likely class after masking in Eq. 2:
$\mathcal{L}_{align}(x_{i},y^{\prime}_{i})=-s_{i,ind(i)}\log p(class_{ind(i)}|x_{i}),$ (7)
where ${ind(i)}$ is the index of the most likely class:
$ind(i)=\arg\max y^{\prime}_{i}.$ (8)
$y^{\prime}_{i}$ is the masked probability distribution for all classes in Eq. 2, $s_{i,ind(i)}$ is its generation probability in Eq. 4 and $class_{ind(i)}$ is the corresponding tokens.
For other tasks, we use NLL loss to maximize of generation probability of $y^{\prime}_{i}$ for CQA and CQG, summarization and KGC:
$\mathcal{L}_{align}(x_{i},y^{\prime}_{i})=-\sum_{j=1}^{|y^{\prime}_{i}|}\log p(y^{\prime}_{i,j}|x_{i},y^{\prime}_{i,<j}).$ (9)
4 Experiments
We evaluate ZOE on three categories of NLP tasks based on changes in conveyed information from input to output: language understanding tasks, language compression tasks and language creation tasks Deng et al. (2021). Language understanding tasks (e.g., NLI, CQA) aim to comprehend and interpret natural language input given a conversation or document context. For a compression task (e.g., summarization), the goal is to concisely describe the most important information in the input (e.g., a document). A creation task (e.g., CQG, KGC) generates output that adds new information on top of input (e.g., dialogue history).
4.1 Datasets
We conduct experiments on eight widely used benchmark datasets: CANARD, CoQAR, CNN/DM, Newsroom, Doc2dial, Mutual, SNLI and QNLI, covering five NLP tasks: CQA, CQG, summarization, KGC and NLI.
  • CANARD (Elgohary et al., 2019) is a benchmark dataset for CQA and CQG. It is built based on QuAC (Choi et al., 2018) and consists of 40k questions with different context lengths, where answers are selected spans from a given section in a Wikipedia article.
  • CoQAR (Brabant et al., 2022) is a large-scale dataset for CQA and CQG. It annotates 53k questions based on CoQA (Reddy et al., 2019), where the documents are from seven diverse domains.
  • CNN/DM (Nallapati et al., 2016) is a well-known summarization dataset, which consists of 313k articles from CNN and Daily Mail. The summary is written by human experts and shown as bullet points. We use the non-anonymized version See et al. (2017).
  • Newsroom (Grusky et al., 2018) is a large-scale summarization dataset which contains 1.3 million articles and expert-written summaries with high diversity.
  • Doc2dial (Feng et al., 2020) is a document-grounded dialogue dataset with 4,800 annotated conversations and an average of 14 turns. Compared to the prior document-grounded dialogue datasets, this dataset covers a variety of dialogue scenes in information-seeking conversations.
  • Mutual (Cui et al., 2020) is a multi-turn reasoning dialogue, consisting of 8,860 manually annotated dialogues based on Chinese student English listening comprehension exams. It is challenging since it requires a model to handle various reasoning problems.
  • SNLI (Bowman et al., 2015) is a large-scale natural language inference benchmark with 570k utterance pairs. Each pair is manually labeled as entailment, neutral, or contradiction with several annotators.
  • QNLI (Wang et al., 2018) is a natural language inference dataset derived from the Stanford Question Answering Dataset v1.1. An utterance is extracted from the passage and paired with the question. Each pair is then manually labeled according to whether the utterance contains the answer to the question.
4.2 Bias types and Dataset Splitting
In this work, we focus on mitigating three types of widely addressed position bias: lead bias (Kedzie et al., 2018), relative position bias (Shinoda et al., 2022) and lexical bias (Gururangan et al., 2018).
  • Lead bias in summarization is that the generated summary is highly correlated to utterances appearing at the beginning of the document (Kedzie et al., 2018).
  • Relative position bias in QA is a phenomenon that a QA model tends to degrade the performance on samples where answers are located in relative positions unseen during training (Shinoda et al., 2022). The relative position is defined as the relative position between utterances containing the last turn answer span and the current turn answer span.
  • Lexical bias is the phenomenon that deep learning models achieve high accuracy by exploiting trigger words or word overlapping (Gururangan et al., 2018; Poliak et al., 2018).
Following previous works (Ko et al., 2020; Shinoda et al., 2022), we split the samples into biased samples and non-biased samples for simulation depending on the bias type in each dataset. In CQG and CQA datasets (CANARD and CoQAR), we select the samples with relative position equaling 0 or 1 as biased samples and the left samples as non-biased samples. In KGC and summarization datasets (Doc2dial, Mutual, CNN/DM and Newsroom), we filter samples where the reference response is highly correlated to the beginning utterance of the given document into biased samples and the left samples into non-biased samples. For NLI datasets (SNLI and QNLI), samples with specific words are filtered into biased samples and other samples are filtered into non-biased samples. The dataset statistics are shown in Table 1.
Table 1: Dataset statistics after splitting.
Dataset Biased
Non-biased
#Train
#Dev
#Test
#Test
CANARD
500
250
3,460
2,440
CoQAR
500
250
3,222
4,873
CNN/DM
500
250
1,421
5,000
Newsroom
500
200
5,000
5,000
Doc2dial
500
250
2,000
5,000
Mutual
500
250
2,000
5,000
SNLI
500
250
2,000
5,000
QNLI
500
250
2,000
5,000
4.3 Evaluation Metrics
Following previous works (Chen et al., 2019; Nallapati et al., 2016; Tuan et al., 2020; Meng et al., 2020), we adopt ROUGE-L Lin (2004) as evaluation metrics for CQA, CQG, summarization and KGC tasks, in which ROUGE-L has been shown to correlate well with human evaluation (Liu and Liu, 2008). We use macro-accuracy for NLI task as NLI is indeed a classification task. We use nlg-eval package [^3] https://github.com/Maluuba/nlg-eval for the implementation of evaluation metrics.
4.4 Baseline Methods
  • Random Position (RT) (Shinoda et al., 2022) randomly perturbs input positions to reduce the model’s dependence on token positions in prediction.
  • FlanT5 (FT) is the FlanT5-large model (Chung et al., 2022) fine-tuned on the dataset for the target task to improve the performance of the target task.
  • MarCQAp (Gekhman et al., 2023) is a novel prompt-based history modeling approach for CQA and CQG that highlights answers from previous conversation turns by inserting textual prompts in their respective positions.
  • Minimax (Korakakis and Vlachos, 2023) leverages an auxiliary model to maximize the learner’s loss by up-weighting ‘hard’ samples, thus reducing the reliance of NLI models on shortcuts in ‘easy’ samples.
  • GenX (Varab and Xu, 2023) is a new summarization paradigm that unifies extractive and abstractive summarization with generative modeling.
  • SG-CQG (Do et al., 2023) is a state-of-the-art CQG models with two stages: what-to-ask for rational span selection in the referential document and how-to-ask for question generation.
  • FocusL (Deng et al., 2023) is a debiasing method built for KGC by adaptively re-weighting the loss of each token, thus encouraging the model to pay special attention to knowledge utilization.
4.5 Implementation Details
We use FlanT5-large (Chung et al., 2022) as the base model for all models. The hidden size is 768. We use the Adam optimizer with a default learning rate 1e-4 (Kingma and Ba, 2015) and set gradient clipping with a default maximum gradient norm of 1.0444https://huggingface.co/docs/transformers/v4.33.0/en/main_classes/trainer. We select the best model based on the BLEU@2 or macro-accuracy score on the validation set. We set $\alpha$=0.1 for CQG and CQA and summarization tasks, and $\alpha$=0.2 for NLI and KGC tasks, by default. To facilitate the reproducibility of the results, we will share the code of all methods and datasets soon.
5 Results
Table 2: Overall performance (%) on language understanding tasks. Boldface indicates the best results in terms of the corresponding metrics.
Method NLI (%) CQA (%)
SNLI QNLI CoQAR CANARD
Biased Non-Biased Biased Non-Biased Biased Non-Biased Biased Non-Biased
T5 77.5 79.0 90.3 88.3 47.4 39.8 34.0 17.3
RP 61.5 51.6 60.4 20.4
MarCQAp 66.9 52.3 66.3 21.3
Minimax 91.6 87.0 90.9 89.7
FT 91.8 86.6 93.0 89.8 64.6 51.9 67.5 20.8
ZOE 91.8 89.0 93.3 91.3 66.3 53.7 65.7 21.9
Table 3: Overall performance (%) on language creation tasks. Boldface indicates the best results in terms of the corresponding metrics.
Method CQG (%) KGC (%)
CoQAR CANARD Doc2dial Mutual
Biased Non-Biased Biased Non-Biased Biased Non-Biased Biased Non-Biased
T5 16.0 17.0 17.9 17.1 23.9 13.7 23.9 21.6
RP 23.2 18.3 24.6 21.2 35.3 32.2 55.3 51.6
MarCQAp 19.7 17.9 26.0 21.8
SG-CQG 14.9 15.5 19.7 17.8
FocusL 39.8 35.9 53.4 52.7
FT 26.7 17.2 26.0 21.6 38.7 36.0 58.8 54.7
ZOE 26.7 18.8 25.9 22.4 40.6 38.3 57.4 56.3
Table 4: Overall performance (%) on language compression tasks. Boldface indicates the best results in terms of the corresponding metrics.
Method Summarization (%)
CNN/DM Newsroom
Biased Non-Biased Biased Non-Biased
T5 22.3 11.6 35.1 20.0
RP 23.9 15.1 47.5 22.0
GenX 17.6 13.7 29.2 19.0
FT 26.9 16.8 51.1 19.8
ZOE 27.1 17.3 50.9 21.3
The overall performances of all methods on language understanding tasks, language creation tasks and language compression tasks are listed in Table 2, Table 3 and Table 4. We have three main observations from the results.
First, LLMs are susceptible to the bias in the dataset after fine-tuning. As we can see in Table 2, after fine-tuning, FT achieves 34.7% improvement on biased samples of CoQAR, but 8.6% improvement on non-biased samples. This is because LLMs can easily overfit the shortcut of the training dataset in fine-tuning, just like existing neural networks (Ko et al., 2020).
Second, ZOE can mitigate position bias significantly on almost all datasets for three types of tasks. As shown in Table 2, Table 3 and Table 4, ZOE improves the performance on non-biased samples, ranging from 1% to 2% on almost all tasks, compared to T5 and all baselines. The reason is that ZOE can leverage pre-trained T5 with lower position bias as a slave objective by MSA and multi-objective optimization.
Third, ZOE only sacrifices a small performance on biased samples when mitigating position bias. As shown in Table 4, RP achieves comparable performance to ZOE on non-biased samples of Newsroom. However, the ROUGE-L of RP drops 3.6% compared to FT on biased samples, while that of ZOE only drops 0.2%. This is because the perturbation in RP impairs the overall data quality for fine-tuning, while ZOE only prune partial objective for optimization in §3.3.
Note that some baselines achieve poor performance, sometimes even lower than T5. First, in Table 4, GenX performs worse even than T5 on biased samples of Newsroom. The reason is that the summarization datasets are abstractive and suitable for generative models, e.g., T5, while GenX (Varab and Xu, 2023) is an extractive baseline. Second, in Table 3, the performance of SG-CQG is worse than that of FT on biased samples of CoQAR. This is because SG-CQG is a three-stage question generation model focusing on improving question diversity. The selected answer span for question generation is randomly chosen from massive generated candidates.
6 Analysis
6.1 Analysis of Unsupervised Responses
[…]
Method CoQAR CANARD
Biased N-Biased Biased N-Biased
ZOE 26.7 18.8 25.9 22.4
- w/o MSA 25.8 18.0 26.1 21.5
- w/ T5-base 26.3 18.3 26.1 22.0
- w/ T5-xlarge 26.6 17.9 25.7 22.0
To verify the effect of the unsupervised responses, we conduct analyses of unsupervised responses in various sources with different qualities in Table 5. ZOE w/o MSA, ZOE w/ T5-base and ZOE w/ T5-xlarge denote ZOE using unsupervised responses without alignment, responses from T5-base and T5-xlarge, respectively. We have two observations.
First, lower response quality leads to worse performance. As we can see, with MSA to improve the quality of unsupervised responses, ZOE outperforms ZOE w/o MSA on non-biased samples. Poor-quality responses undermine the model comprehension of the task, thus leading to worse performance.
Second, using various sources of unsupervised responses may degrade model performance. ZOE w/ T5-base and ZOE w/ T5-xlarge leverage unsupervised responses from T5-base and T5-xlarge, while they both perform worse than ZOE on CANARD. We infer that responses from T5-base or T5-xlarge use different knowledge/parameters to generate questions, which mismatch with that of T5. This difference amplifies the divergence between master objective and slave objective in §3.2. Even so, ZOE still outperforms the fine-tuned T5.
6.2 Analysis of Objective Weighting
[CQA_CANARD]Refer to caption [CQG_CoQAR]Refer to caption
[Newsroom]Refer to caption [Doc2dial]Refer to caption
Figure 3: Performance (%) over each $\alpha$ of CQA on CANARD, CQG on CoQAR, summarization on Newsroom, KGC on Doc2dial. The x-axis denotes the value of $\alpha$ and the y-axis denotes the ROUGE-L on non-biased samples.
To analyze the effect of weighting on slave objective, we present the performance of ZOE using different $\alpha$ in Fig. 3. FT denotes the fine-tuned LLM. We have two observations from the results.
First, the performance of ZOE drops with the increase of the weight of unsupervised responses in multi-objective optimization. In CQA on CANARD, the ROUGE-L score of ZOE drops from 21.9% to 21.5% when ranging $\alpha$ from 0.1 to 0.5. This is because the performance on non-biased samples depends on not only the degree of bias but also the quality on training samples. Increasing $\alpha$ will reduce the position bias of all responses, yet it will hurt the quality concurrently.
Second, our proposed ZOE outperforms fine-tuned T5 under various $\alpha$. As shown in Fig. 3, the CQA performances of ZOE on CANARD all exceed 21.5% using various $\alpha$, while FT only achieves ROUGE-L value 20.8%. This demonstrates the effectiveness and robustness of ZOE in mitigating position bias.
6.3 Analysis on Training Samples
[CQA_CANARD]Refer to caption [CQG_CoQAR]Refer to caption
[Newsroom]Refer to caption [Doc2dial]Refer to caption
Figure 4: Performance (%) over different numbers of training samples of CQA on CANARD, CQG on CoQAR, summarization on Newsroom, KGC on Doc2dial. The x-axis denotes the number of training samples and the y-axis denotes the ROUGE-L on non-biased samples dataset.
We also analyze the effect of the number of training samples to verify the effectiveness of ZOE under various low-resource settings. We plot the results in Fig. 4 and have two observations from the results.
First, all methods always perform better when increasing training samples. As we can see in Fig. 4, the ROUGE-L score of FT increases from 16.0% to 18.5% on CoQAR when the number of training samples increases from 50 to 1,000. The increase of training samples can improve the generalization ability of LLMs, thus leading to better performance on non-biased samples.
Second, our proposed ZOE outperforms fine-tuned T5 under various settings on training samples. As shown in Fig. 4, the ROUGE-L scores of ZOE depicted by the orange charts are consistently higher than that of FT in blue. This is because there are always around 40% unsupervised responses after aligning for fine-tuning when ranging the number of training samples, which are enough for effectively debiasing.
6.4 Case Study
Table 6: An example of CQA on CANARD. U1, U4, and U5 are the 1st, 4th, 5th utterances in the document and T1 is the first turn utterance of the dialogue. ‘Target’ is the target answer of the document and ‘FT’ and ‘ZOE’ are the generated answers. ‘Position’ is the position of the utterance containing the answer in the document and ‘Relative position’ is the distance of grounded utterances between the current answer and the last answer.
ID
Document
U1
Gaston had worked with players at an individual level as a hitting instructor …
U4
In the six games the Blue Jays played in those places during World Series play
U5
… and Gaston was the first ever African-American manager to win a World Series.
ID
Context
Position
T1
Question: What happens in the series?
Answer: In the six games the Blue Jays played in those places during World Series play, the Jays went 4-2 …
U4
T2
Question: What else did Cito do?
Model
Question
Relative position
Target
Gaston had worked with players at an individual level
-3 (U1)
FT
Gaston was the first ever African-American manager to win a World Series.
1 (U5)
ZOE
Gaston had worked with players at an individual level
-3 (U1)
Table 7: An example of CQG on CANARD. U1 - U5 are the utterances in the document and T1 is the first turn utterance of the dialogue. ‘Target’ is the target question of the document, ‘T5-1’ and ‘T5-2’ are the questions generated by pre-trained T5. ‘Position’ is the position of the utterance containing the answer to the question in the document and ‘Relative position’ is the distance of grounded utterances between the current question and the last question.
ID
Document
U1
Gautam Gambhir, born 14 October 1981, is an Indian cricketer, …
U2
Gambhir was picked up by the Delhi Daredevils franchise in the first player auction of the Indian Premier League for a price of US\$725,000 a year.
U3
He became the second highest run-scorer of the inaugural season with …
U4
He was promoted to the post of Captain of the Delhi Daredevils for IPL Season 2010.
U5
At the end of the tournament he became the only player from Delhi Daredevils to score more than 1000 runs in the IPL.
U6
In the 2011 IPL player auction, …
ID
Context
Position
T1
Question: Who was Gautam Gambhir in Indian premier league?
Answer: He was promoted to the post of Captain of the Delhi Daredevils for IPL Season 2010.
U4
Source
Question
Relative position
Target
What did Gautam Gambhir do as captain?
1 (U5)
T5-1
Where was Gautam Gambhir born?
-3 (U1)
T5-2
Was Gautam Gambhir in Indian premier league?
-2 (U2)
To investigate the reason for ZOE’s effectiveness, we present an example of generated responses in CQA in Table 6 and an example of unsupervised responses in CQG in Table 7.
In CQA example, ZOE can generate an answer with lower bias than FT. As shown in Table 6, FT generates an answer from U5, adjacent to U4, the utterance containing the last turn answer. In contrast, ZOE generates an answer from U1, which is far from U4. After fine-tuning with unsupervised responses from various positions, ZOE cannot easily fall into the trap of finding answers from neighboring utterances around the utterance containing the last turn answer.
In CQG example in Table 7, ZOE uses unsupervised responses which have lower bias than the target response. As we can see, target question is based on the utterance with relative position as 1, while unsupervised questions generated from T5 are based on utterances varying in the document. Fine-tuning with T5-generated unsupervised responses encourages the model to generate questions based on utterances in various document positions, thus mitigating position bias.
7 Related Work
7.1 Position Bias and Debiasing
Position bias has been explored mainly in three extractive tasks, i.e., NLI, summarization and CQA. Besides, there are also other explored extractive tasks but with relatively less attention, e.g., sentiment analysis (Ma et al., 2021). In, NLI, Gururangan et al. (2018); Poliak et al. (2018) show that the class labels are highly correlated to certain words in the hypothesis. McCoy et al. (2019) report that models always rely on word overlap between hypothesis and premise for prediction. Karimi Mahabadi et al. (2020); Du et al. (2021) train a robust NLI model in an ensemble manner by PoE. They train a hypothesis model to learn the lexical bias, which guides the NLI model to focus on other patterns in the dataset that generalize better.
There are similar findings on position bias in CQA tasks. Weissenborn et al. (2017); Sugawara et al. (2018) demonstrate that only using partial inputs is sufficient to correctly extract the answer span for the question in most cases. Ko et al. (2020) address that the absolute position of answer spans can work as a spurious bias as models trained on biased datasets fail to generalize over unseen answer positions. Shinoda et al. (2022) report that the relative position of answer spans is another clue for position bias. To mitigate position bias, Ko et al. (2020); Shinoda et al. (2022) build bias ensemble models by PoE similar to (Karimi Mahabadi et al., 2020), where they design biased models based on position-only features to guide the CQA models to leverage more semantic features for answering.
Kedzie et al. (2018); Grenander et al. (2019) find 58% of selected summary utterances come directly from the lead utterances, and models trained on these articles perform considerably worse when utterances in the article are randomly shuffled. To mitigate lead bias, a simple but effective method is to randomly shuffle the document for summarization training (Grenander et al., 2019). Then, Xing et al. (2021) uses adversarial training for debiasing. They design a position prediction module and optimize the reverse loss for position prediction, forcing the encoder to leverage non-position features. Besides, Zhu et al. (2021) leverage lead bias for unsupervised training, i.e., treating the leading utterances from a news article as the target summary and the rest as the article.
However, existing works on mitigating position bias always focus on extractive tasks. They can hardly be transferred to generative tasks as the label space in generative tasks is too large for bias estimation or adversarial training. In this paper, we focus on mitigating position bias for generative LLMs.
7.2 Bias in LLMs and Debiasing
Works on debiasing for LLMs always focus on social bias, e.g., gender bias and racial bias, rather than position bias (Meade et al., 2022; Du et al., 2022). Existing works on mitigating social bias for LLMs can be classified into three types: pre-processing methods, in-processing methods and post-processing methods.
In pre-processing, Zmigrod et al. (2019) adopt a counterfactual data augmentation (CDA) algorithm to mitigate social bias by swapping bias attribute words (e.g., he/she) in a dataset. Choi et al. (2022) modify CDA by masking the terms casual to label to force the model to learn label-invariant features.
In in-processing, Guo et al. (2022); Li et al. (2023); Yang et al. (2023) propose a two-stage adversarial method for debiasing, where they first train a continuous prompt to enlarge the bias of utterance pairs and then force the LLMs to minimize the difference of utterance pairs using the prompt.
In post-processing, Schick et al. (2021) propose a decoding algorithm that reduces the probability of a model producing biased text using a textual description of the undesired behavior for prompting. Meade et al. (2023) propose an ICL strategy which leverages non-biased demonstrations to guide the generation on safety.
However, existing methods for mitigating social bias either require the external bias knowledge for data augmentation and training adversarial prompts, or require a non-biased dataset for building demonstrations, which is unpractical in application. Differently, we propose a zero-shot framework for LLMs on mitigating position bias, without knowing any external bias knowledge or non-biased samples, which is general, simple but effective.
8 Conclusion
In this paper, we have proposed a zero-shot debiasing ZOE framework for LLMs. We adopt multi-objective optimization to mitigate position bias for LLMs, with a slave objective based on unsupervised responses. These responses are aligned by a proposed MSA module based on the target responses. Extensive experiments on five tasks and eight benchmark datasets show that ZOE outperforms existing baselines on non-biased samples while retaining performance or sacrificing little performance on biased samples. It proves that leveraging aligned unsupervised responses is a practicable solution to mitigate position bias for generative LLMs.
9 Limitations
This work has the following limitations. First, ZOE needs a pre-trained LLM to generate unsupervised responses with low bias, where these responses are still biased. Second, the final performance of ZOE depends on the quality of unsupervised responses, which are still noisy after the proposed MSA. In future work, we plan to address these issues by investigating non-biased models from other domains and model-based strategies to align unsupervised responses.
Ethical Considerations
We realize that there are risks in developing dialogue systems, so it is necessary to pay attention to the ethical issues of dialogue systems. We have used the public pre-trained LLMs, e.g., FlanT5-large, and public datasets, e.g., CANARD and CoQAR, to conduct the experiments. All models and datasets are carefully processed by publishers to ensure that there are no ethical problems.
Appendix A Analysis
Table 8: Overall performance (%) on language understanding tasks. Boldface indicates the best results in terms of the corresponding metrics.
Method NLI (%) CQA (%)
SNLI QNLI CoQAR CANARD
Biased Non-Biased Biased Non-Biased Biased Non-Biased Biased Non-Biased
ZOE 91.8 89.0 93.3 91.3 66.3 53.7 65.7 21.9
- w/o MSA 91.7 88.4 93.7 90.1 66.3 53.7 65.7 21.9
- w/ T5-base 91.7 87.7 93.1 91.0 64.3 52.3 61.6 19.9
- w/ T5-xlarge 92.2 88.9 93.9 91.2 65.1 53.6 67.1 22.4
Table 9: Overall performance (%) on language creation tasks. Boldface indicates the best results in terms of the corresponding metrics.
Method CQG (%) KGC (%)
CoQAR CANARD Doc2dial Mutual
Biased Non-Biased Biased Non-Biased Biased Non-Biased Biased Non-Biased
ZOE 26.7 18.8 25.9 22.4 40.6 38.3 57.4 56.3
- w/o MSA 25.8 18.0 26.1 21.5 32.6 34.1 55.9 53.5
- w/ T5-base 26.3 18.3 26.1 22.0 42.8 36.3 60.6 55.4
- w/ T5-xlarge 26.6 17.9 25.7 22.0 44.1 39.1 58.7 56.9
Table 10: Overall performance (%) on language compression tasks. ‘N-Biased’ denotes performance on non-biased samples. Boldface indicates the best results in terms of the corresponding metrics.
Method Summarization (%)
CNN/DM Newsroom
Biased N-Biased Biased N-Biased
ZOE 27.1 17.3 50.9 21.3
- w/o MSA 27.1 14.8 48.0 20.2
- w/ T5-base 26.8 14.2 47.7 20.3
- w/ T5-xlarge 28.0 14.6 48.4 20.2
A.1 Analyses of Unsupervised Responses
Table 8, Table 9 and Table 10 demonstrate ZOE performance with different qualities of unsupervised responses. As we can see, lower response quality leads to worse performance. Besides, using various sources of unsupervised responses may improve or degrade model performance depending on the task type.
A.2 Analyses of Objective Weighting
Fig. 5 provides the overall performance on different $\alpha$. In most cases, ZOE outperforms fine-tuned T5.
A.3 Analyses on Training Samples
Fig. 6 provides the overall performance on different numbers of training samples. In most cases, ZOE outperforms fine-tuned T5.
[CQG_CANARD]Refer to caption [CQG_CoQAR]Refer to caption
[CQA_CANARD]Refer to caption [CQA_CoQAR]Refer to caption
[Newsroom]Refer to caption [CNN/DM]Refer to caption
[Doc2dial]Refer to caption [Mutual]Refer to caption
[SNLI]Refer to caption [QNLI]Refer to caption
Figure 5: Performance over each $\alpha$ of CQG on CANARD, CQG on CoQAR, CQA on CANARD, CQA on CoQAR, summarization on Newsroom, summarization on CNN/DM, KGC on Doc2dial, KGC on Mutual, NLI on SNLI, NLI on QNLI. The x-axis denotes the value of $\alpha$ and the y-axis denotes the ROUGE-L on non-biased samples.
[CQG_CANARD]Refer to caption [CQG_CoQAR]Refer to caption
[CQA_CANARD]Refer to caption [CQA_CoQAR]Refer to caption
[Newsroom]Refer to caption [CNN/DM]Refer to caption
[Doc2dial]Refer to caption [Mutual]Refer to caption
[SNLI]Refer to caption [QNLI]Refer to caption
Figure 6: Performance over different numbers of training samples of CQG on CANARD, CQG on CoQAR, CQA on CANARD, CQA on CoQAR, summarization on Newsroom, summarization on CNN/DM, KGC on Doc2dial, KGC on Mutual, NLI on SNLI, NLI on QNLI. The x-axis denotes the number of training samples and the y-axis denotes the ROUGE-L on non-biased samples dataset.